OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\sqrt{98^5}\]

OpenStudy (anonymous):

Need to learn the steps properly :P

OpenStudy (uri):

Step 1:Logout Step 2:Go to a stationery Step 3:Buy a scientific calculator.

OpenStudy (anonymous):

...

OpenStudy (anonymous):

*goes to laptop scientific calculator*

OpenStudy (anonymous):

The choices are like: \[2x^2 \sqrt{7x}\]

OpenStudy (anonymous):

-.- Help me please.

OpenStudy (luigi0210):

\[\sqrt{(2*49)^5}\]

OpenStudy (anonymous):

What would I do next? I don't understand .-.

OpenStudy (anonymous):

why dont you first rewrite it like this \[\sqrt{98^4*98}\]

OpenStudy (anonymous):

Then what would I do? I don't see what the next step would be to do.

OpenStudy (anonymous):

what would happen if we put out the 98^4

OpenStudy (luigi0210):

Oh i see. Think of it this way: \[98^{4/2}*\sqrt{49*2}\]

OpenStudy (anonymous):

Divisible by 2?

OpenStudy (anonymous):

so yeah you would actually get \[98^2*\sqrt{98}\]

OpenStudy (luigi0210):

Or simplified down more would be: \[7*98^2 \sqrt{2}\]

OpenStudy (anonymous):

How do you get it to 2?

OpenStudy (luigi0210):

Breaking it down.

OpenStudy (anonymous):

Dividing 98 by 49?

OpenStudy (anonymous):

yeah you can rewrite 98 as 49*2

OpenStudy (luigi0210):

Yup, yup

OpenStudy (anonymous):

Oh okay. Well I have a few more questions could one of you help? They look easier than this one xD

OpenStudy (luigi0210):

And the perfect square of 49 is 7

OpenStudy (anonymous):

Oh that's how you got 7

OpenStudy (luigi0210):

Yup :D

OpenStudy (anonymous):

Makes more sense now :P

OpenStudy (anonymous):

a lot more sense xDD

OpenStudy (anonymous):

With the information I was given the answer is: \[7x^2 \sqrt{2}\] correct?

OpenStudy (anonymous):

I wrote the first thing wrong. \[\sqrt{98x^5}\]

OpenStudy (anonymous):

you need that to be simplified?

OpenStudy (luigi0210):

Yea thats right

OpenStudy (luigi0210):

But you're missing a piece

OpenStudy (anonymous):

Which is?o;

OpenStudy (luigi0210):

\[\sqrt{49*2*x^4*x}\]

OpenStudy (anonymous):

So it's \[7x^2\sqrt{2x}\]

OpenStudy (luigi0210):

There you go :)

OpenStudy (anonymous):

Can you help me with a few more?o;

OpenStudy (anonymous):

\[\sqrt{48}\] is my next one.

OpenStudy (luigi0210):

Well try them first then post a new question if you need help

OpenStudy (anonymous):

Alrighty. (:

OpenStudy (luigi0210):

\[\sqrt48=\sqrt{4*4*3}\]

OpenStudy (anonymous):

I don't know how to do them at all, that's why I asked for steps to learn :P

OpenStudy (anonymous):

\[4 \sqrt{3}\] is that equal to \[\sqrt{4*4*3}\]

OpenStudy (luigi0210):

Yea :)

OpenStudy (anonymous):

Can you explain why? :P

OpenStudy (luigi0210):

I'm awful at explanations.. sorry brah D:

OpenStudy (anonymous):

How do you know it's equal to eachother then? :P

OpenStudy (anonymous):

because \[\sqrt{4*4*3}=\sqrt{4^2*3}\] so since its 4^2 when we pull it out of the squareroot we look for \[\sqrt{4^2}=4\] so we have \[4*\sqrt{3}\] left

OpenStudy (anonymous):

Understandable :P

OpenStudy (anonymous):

Let me get your opinion. on this.

OpenStudy (anonymous):

\[\sqrt{100}\] would be equal to \[50\sqrt{2}\] ?

OpenStudy (anonymous):

not necessarily :p

OpenStudy (anonymous):

absolutely not xD

OpenStudy (anonymous):

Dang.-.

OpenStudy (anonymous):

Ok so u'd find the factors of 100 which are: 10 * 10. Since it's got 2 of the same factor u can take it out of the square root